OpenStudy (christos):

Logarithmic differantiation, y = (x^3 -2x)^ln(x) I got a very long result I dont know if my solution is correct

OpenStudy (christos):

y'=(x^3 - 2x)^ln(x)lnx^3-2x+lnx*(2x-2)/(x^3-2x)

OpenStudy (christos):

I forgot an 1/x next to (x^3 - 2x)

OpenStudy (christos):

multiplication

sam (.sam.):

I got\[\Large y'=[\frac{(3x^2-2)\ln(x)}{x^3-2x}+\frac{\ln(x^3-2x)}{x}](x^3-2x)^{\ln(x)}\]

OpenStudy (christos):

hm

OpenStudy (reemii):

only thing to correct in your answer: it's \(3x^2-2\) instead of \(2x-2\), and you forgot some parentheses.

OpenStudy (christos):

Yea I just noticed that error, and the parenthesis I know, I just surpassed them. Other than that all ok ??

OpenStudy (reemii):

yes

sam (.sam.):

Yeah \[y=(x^3-2x)^{\ln(x)} \\ \\ \ln(y)=\ln(x^3-2x)^{\ln(x)} \\ \\ \ln(y)=\ln(x)\ln(x^3-2x)\] Implicit differentiation , and product rule + chain rule for RHS \[\frac{1}{y}y'=\ln(x)\frac{1}{x^3-2x}(3x^2-2)+\ln(x^3-2x)(\frac{1}{x})\]

OpenStudy (christos):

thank you guys